zarasht:

A solution is made using 154.7 g of toluene (MM = 92.13 g/mol) and 80.0 g of benzene (MM = 78.11 g/mol). What is the mole fraction of the toluene in the solution?

justjm:

The formula for mole fraction is really just \(X_n = \frac{\text{moles of n}}{\text{moles of total solution}}\) Here you can use dimensional analysis to help you get from grams to a ratio It is asking you to find mole fraction of toluene, so find the mols of toluene first \(154.6\ g\ \cdot\frac{1\ mol\ }{92.13\ g}=\text{mol toluene\\}\) Now to find the mols of the whole solution, you need to add the mols of toluene AND the mols of benzene...dimensional analysis needed again \(80.0\ g\ \cdot\frac{1\ mol}{78.11\ g}= \text{mol Benzene}\) Once you find those values Divide the moles of them together by the mols of toluene to get the mol fraction, X

zarasht:

im a little bit confused

justjm:

Okay, I'll get you through a few more steps. Using the mol equations we set up, we find that there are 154.6/92.13 = 1.69 mols of Toluene and there are 80.0g/78.11 = 1.02 mols of Benzene Now to find the mole fraction for toluene, or \(X_{toluene}\), you would do \(\frac{1.69~mols~Toluene}{1.69~+~1.02}\) Get it better?

zarasht:

yes thanks

justjm:

np